How many liters of a 9% solution of salt should be added to a 27% solution in order to obtain 1062 liters of a 15% solution?
Answer by ankor@dixie-net.com(22740)

How many liters of a 9% solution of salt should be added to a 27% solution in order to obtain 1062 liters of a 15% solution?
:
let x = amt of 27% required
then
(1062-x) = amt of 9% required
:
A typical mixture equation
.27x + .09(1062-x) = .15(1062)
.27x + 95.58 - .09x = 159.3
.27x - .09x = 159.3 - 95.58
.18x = 63.72
x = 63.72/.18
x = 354 liters of 27% solution
then
1062 - 354 = 708 liters of 9% needs to added to the above solution